OpenStudy (anonymous):

I need someone to explain how to do this A $6000 investment earns $690 simple interest in two years. What is the annual interest rate? A. 1.15% B. 5.75% C. 11.5% D. 57.5%

OpenStudy (okdutchman7):

First divide $690 by 2 because the interest rate has been compounded over 2 years.

OpenStudy (okdutchman7):

@LoveIt

OpenStudy (anonymous):

345 @okdutchman7 Wouldn't I multiply them though?

OpenStudy (okdutchman7):

No because interest is compounded on a yearly basis

OpenStudy (okdutchman7):

The next step is to divide 345 by 6000.

OpenStudy (okdutchman7):

^This is the amount of interest that is received each year.

OpenStudy (okdutchman7):

@LoveIt Let me know when you are done with that.

OpenStudy (anonymous):

o hold on

rebeccaxhawaii (rebeccaxhawaii):

mm

OpenStudy (anonymous):

17.3913...

OpenStudy (anonymous):

its 5.75%

OpenStudy (anonymous):

just solved it

OpenStudy (okdutchman7):

@Joshtray. No direct answers.

OpenStudy (anonymous):

oh sorry, 0.0575

OpenStudy (okdutchman7):

It can't be 17.3913... because 345 is smaller than 6000. Try it again.

OpenStudy (okdutchman7):

That's right and when multiplied by 100 leaves you with 5.75%

OpenStudy (okdutchman7):

Good Job! @LoveIt

OpenStudy (anonymous):

Thanks! I kind of just realized it said ANNUAL interest so now it makes sense, heh

OpenStudy (okdutchman7):

I'm glad that you understand, and I was happy to help anyway.

OpenStudy (anonymous):

Rate = 100 x Interest\[Rate = \frac{100 \times Interest }{ Principle \times Time }\]

OpenStudy (anonymous):

Thanks Joshtray!